Solve the following partial fractions
(i) 3x-1/2x(x^2 +2)

(ii) 2x+1/(x-1)(x^2 +1)

(iii) x^3 +x+1/x(x^2 -x+1)

(iv) 3x+2/(x-2)(x^2 +3x+1)
1
Expert's answer
2020-06-30T17:16:10-0400

(i)


"{3x-1 \\over 2x(x^2+2)}={A \\over x}+{Bx \\over x^2+2}+{C \\over x^2+2}\t\n="

"={Ax^2+2A+Bx^2+Cx\over 2x(x^2+2)}"

"A+B=0"

"C=\dfrac{3}{2}"

"2A=-\dfrac{1}{2}"

"A=-\dfrac{1}{4}, B=\dfrac{1}{4}, C=\dfrac{3}{2}"


"{3x-1 \over 2x(x^2+2)}={-\dfrac{1}{4} \over x}+{\dfrac{1}{4}x \over x^2+2}+{\dfrac{3}{2}\over x^2+2}"

(ii)


"{2x+1 \\over (x-1)(x^2+1)}={A \\over x-1}+{Bx \\over x^2+1}+{C \\over x^2+1}\t\n="

"={Ax^2+A+Bx^2-Bx+Cx-C \over (x-1)(x^2+1)}"

"A+B=0"

"-B+C=2"

"A-C=1"

"A=\dfrac{3}{2}, B=-\dfrac{3}{2}, C=\dfrac{1}{2}"


"{2x+1 \over (x-1)(x^2+1)}={\dfrac{3}{2} \over x-1}+{-\dfrac{3}{2}x \over x^2+1}+{\dfrac{1}{2} \over x^2+1}"

(iii)


"{x^3+x+1 \over x(x^2-x+1)}=1+{x^2+1 \over x(x^2-x+1)}="

"=1+{A \over x}+{Bx \over x^2-x+1}+{C \over x^2-x+1}="

"=1+{Ax^2-Ax+A+Bx^2+Cx \over x(x^2-x+1)}"

"A+B=1"

"-A+C=0"

"A=1"

"A=1,B=0,C=1"


"{x^3+x+1 \over x(x^2-x+1)}=1+{1\over x}+{ 1\over x^2-x+1}"

(iv)


"{3x+2 \over (x-2)(x^2+3x+1)}"


Let "x^2+3x+1=0"

"D=(3)^2-4(1)(1)=5"

"x=\dfrac{-3\pm\sqrt{5}}{2}"


"x^2+3x+1=(x-\dfrac{-3-\sqrt{5}}{2})(x-\dfrac{-3+\sqrt{5}}{2})="

"={1\over 4}(2x+3+\sqrt{5})(2x+3-\sqrt{5})"

"{3x+2 \over (x-2)(x^2+3x+1)}={A \over x-2}+{2B \over 2x+3+\sqrt{5}}+{2C \over 2x+3-\sqrt{5}}="

"={4Ax^2+12Ax+4A+(x-2)(4Bx+6B-2\sqrt{5}B+4Cx+6C+2\sqrt{5}C)\over 4(x-2)(x^2+3x+1)}"

"4A+4B+4C=0"

"12A-8B+6B-2\sqrt{5}B-8C+6C+2\sqrt{5}C=3"

"4A-12B+4\sqrt{5}B-12C-4\sqrt{5}C=2"


"A+B+C=0"

"-12B-12C-2B-2\sqrt{5}B-2C+2\sqrt{5}C=3"

"-16B+4\sqrt{5}B-16C-4\sqrt{5}C=2"


"A+B+C=0"

"-14B-14C-2\sqrt{5}B+2\sqrt{5}C=3"

"-8B+2\sqrt{5}B-8C-2\sqrt{5}C=1"


"A=\dfrac{2}{11}"

"B=-\dfrac{1}{11}-\dfrac{\sqrt{5}}{44}"

"C=-\dfrac{1}{11}+\dfrac{\sqrt{5}}{44}"


"{3x+2 \over (x-2)(x^2+3x+1)}={\dfrac{2}{11} \over x-2}+{-\dfrac{2}{11}-\dfrac{\sqrt{5}}{22} \over 2x+3+\sqrt{5}}+{-\dfrac{2}{11}+\dfrac{\sqrt{5}}{22} \over 2x+3-\sqrt{5}}"